OpenStudy (nickszuhy):

Find the x-intercepts of the parabola with vertex (6,27) and y-intercept (0,-81). Write your answer in this form (x1,y1),(x2,y2). If necessary, round to the nearest hundredth

satellite73 (satellite73):

you know how to put this in " vertex form" as \[y=a(x-h)^2+k\]?

OpenStudy (nickszuhy):

It tells me use this formula: (y-k)=a(x-h)^2

satellite73 (satellite73):

ok same thing really

satellite73 (satellite73):

put in \(h=6, k=27\) leave the \(a\) as an \(a\) we will find it in a second

satellite73 (satellite73):

write it, let me check, then we can find \(a\) and finally we can solve

OpenStudy (nickszuhy):

all i have then is: (y-6)=a(x-27)^2

satellite73 (satellite73):

yes, now to find \(a\)

OpenStudy (nickszuhy):

But what do i do with y? that has to have a number for it

OpenStudy (nickszuhy):

And what is x?

satellite73 (satellite73):

oh wait, it is wrong

satellite73 (satellite73):

we got the x and y backwards here \((y-6)=a(x-27)^2 \) it should have been \[(y-27)=a(x-6)^2 \]

OpenStudy (nickszuhy):

I have 2 more questions to answer that i might need help on

satellite73 (satellite73):

we are not done with this one yet now we need to find \(a\)

satellite73 (satellite73):

put \(x=0,y=-81\) and solve \[-81-27=a(0-6)^2\] for (a\)

OpenStudy (nickszuhy):

so then from the y intercept i get (-81-27)=a(0-6)^2

satellite73 (satellite73):

right, solve for \(a\) let me know when you get it

OpenStudy (nickszuhy):

-3=a

satellite73 (satellite73):

ok so now we have our parabola nailed down it is \[y-27=-3(x-6)^2\]

satellite73 (satellite73):

i would prefer it looked like \[y=-3(x-3)^2+27\] so now we can set \(y=0\) and solve for \(x\) to finish the question

satellite73 (satellite73):

ok typo there \[y=-3(x-6)^2+27\]

satellite73 (satellite73):

or just leave it as \(y-27=-3(x-6)^2\) put \(y=0\) solve \[-27=-3(x-6)^2\]

satellite73 (satellite73):

you know how to do that?

OpenStudy (nickszuhy):

so i would get y=-3x^2-6x^2 correct?

satellite73 (satellite73):

no

OpenStudy (nickszuhy):

could you show me

satellite73 (satellite73):

\[-27=-3(x-6)^2\] divide both sides by \(-3\) first then what do you get?

OpenStudy (nickszuhy):

9=(x-6)^2

satellite73 (satellite73):

ok now take the square root of both sides don't forget the \(\pm\) you are going to get two answers

OpenStudy (nickszuhy):

9=x^2*6^2

OpenStudy (nickszuhy):

then divide it all by 9?

satellite73 (satellite73):

hmm o

satellite73 (satellite73):

not square, take the square root

satellite73 (satellite73):

9 has two square roots, a positive one and negative one

OpenStudy (nickszuhy):

yeah 3 and -3

satellite73 (satellite73):

k so from \[9=(x-6)^2\] we get two equations \[x-6=-3\\ x-6=3\]

satellite73 (satellite73):

add 6 to both sides to get your two answers

OpenStudy (nickszuhy):

x=3 x=9

satellite73 (satellite73):

yes, final answers \[(3,0), (9,0)\]

OpenStudy (nickszuhy):

It was right thanks! could you help me out on my other 2 questions that are the same thing.

satellite73 (satellite73):

sure post in a new thread, this one is gettting long